# Cylindrical Gravitational Waves in Einstein-Aether Theory

### 30 Dec 2023 v1

@@ 1 Introduction
Immediately after completing the theory of General Relativity [1] and proposing the three classical tests [2] to verify the theory experimentally, in 1916, Einstein predicted the existence of gravitational waves as a natural consequence of dynamical spacetime [3]. In this first article [3] on gravitational waves, Einstein concluded that accelerated massive objects generate changes in the spacetime curvature, which propagate like waves with the speed of light. However, he concluded that gravitational waves carry no energy because of a calculational error, which he corrected in 1918 [4]. He also derived a famous formula for the energy loss of a system emitting gravitational radiation. Nevertheless, Einstein tried to disprove the existence of gravitational waves in 1937 [5]. The issue of gravitational waves remained controversial until Bondi [6] settled it for good, which was later explored in a series of papers by Bondi and his collaborators. The first experimental evidence, albeit indirect, came in 1979 with the discovery of the orbital decay of the newly discovered binary pulsar by Hulse and Taylor [7]. Finally, the gravitational waves were directly detected in 2015 by the LIGO and Virgo collaborations [8]. Now, with this direct detection of gravitational waves, we have a new window into the universe and we can test our understanding of gravity at higher energies like never before. A fascinating account of the history of gravitational waves can be found in references [9, 10, 11].
@@ 4 Gravitational Wave Equation
Here, we have checked the possible existence of a bifurcation of solutions. We have two possibilities from equation (23): (i) $c_{2}=c_{13}=0$ or (ii) $c_{2}=-c_{13}$. In general we can assume that $H_{0}\neq\Psi_{0}$. In the case (i) we obtain from the equation (16) and from equation (17) the quantitity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(-c_{4}+c_{3})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{3}=c_{4}$ into equations (15)-(20) we obtain exactly the GR field equations. In the case (ii) we obtain again from the equation (16) and from equation (17) the quantitity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(c_{14})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{14}=0$ into equations (15)-(20) we obtain exactly the EA field equations. These two cases suggest no GR limit for EA wave equations.
@@ 5 Solution of the Gravitational Wave Equation
${k}=w\,\sqrt{1-{c_{13}}},$ (29)
where $w$ is the frequency of the gravitational wave. Notice that in EA, the velocity of the wave can be infinity when $c_{13}\rightarrow 1$, while in GR the velocity of the wave is the speed of the light. We also must have $c_{13}<1$ in order to have a wave.
@@
$\Psi{(r,t)}=f\left(r\right)\cos\left(w\,t\right),$ (30)
@@
Let us suppose that $\sqrt{C_{12}}=C_{5}=k$, $C_{8}=C_{6}$, $C_{11}=C_{4}$, $C_{10}=C_{3}$, for the sake of simplicity, and using equation (29), we can rewrite the metric functions as
$\Psi{(r,t)}={C_{1}}\,\cos\left(w\,t\right)\,J_{0}\left(\sqrt{1-{c_{13}}}\,w\,r\right),$ (42)
$H(r,t)$ $=$ $2{C_{8}}\,J_{0}\left(\sqrt{1-{c_{13}}}\,w\,r\right)\left[{C_{3}}\,\cos\left(w\,t\right)+{C_{4}}\,\sin\left(w\,t\right)\right],$
and using equation (41) we have
${\frac{\sqrt{c_{14}}\,\sqrt{1-{c_{13}}}}{\sqrt{{c_{2}}+{c_{13}}}}}=1.$ (44)
@@
$H(r,t)$ $=$ $\cos\left(\,\,w\,t\right)J_{0}\left(\sqrt{1-{c_{13}}}\,w\,r\right).$ (45)
Substituting equation (41) into (45) we get
$H(r,t)$ $=$ $\cos\left(w\,t\right)J_{0}\left(\sqrt{1-{c_{13}}}\,w\,r\right).$ (46)
Comparing equations (42) and (46), we can see that they have the same functional dependence in the coordinates $r$ and $t$. Section 6 will show that this temporal behavior is different in GR.
Figure 1: Wave components $\Psi(w)$ and $H(w)$ using $r=1$, $t=1$ and $c_{13}=0,\pm 3/4$, in equations (42) and (45).
Figure 2: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and for a value of $c_{13}=3/4$.
Figure 3: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and for a value of $c_{13}=0$.
Figure 4: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and for a value of $c_{13}=3/4$.
Figure 5: Wave component pulse $H$ of width $\beta$, using $C_{3}=1$, $C_{4}=0$ and $C_{6}=1/2$, in equation (45) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and for a value of $c_{13}=3/4$.
Figure 6: Wave component pulse $H$ of width $\beta$, using $C_{3}=1$, $C_{4}=0$ and $C_{6}=1/2$, in equation (45) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and for a value of $c_{13}=0$.
Figure 7: Wave component pulse $H$ of width $\beta$, using $C_{3}=1$, $C_{4}=0$ and $C_{6}=1/2$, in equation (45) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (48), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and for a value of $c_{13}=3/4$.
@@ 6 Gravitational Wave Pulse with Width
Let us now calculate the superposition of waves of the type of (42) and (46) with amplitude factor as $\exp(-\beta\,w)$. Here, we will follow the same procedure used in reference [34]. Thus, we have
@@
assuming that $\beta=1$, $t=0$ or $t=2$ or $t=6$ or $t=10$ and $c_{13}=0$ or $c_{13}=\pm 3/4$, we can write that
$\Psi(x,t=0,c_{13}=3/4)$ $=$ $\,{\frac{2}{\sqrt{4+{x}^{2}}}},$ (51)
$\Psi(x,t=2,c_{13}=3/4)=$
$\sqrt{2}{\frac{\sqrt{\sqrt{13120+2496\,x-64\,{x}^{2}-24\,{x}^{3}+{x}^{4}}-140+\left(-6+x\right)^{2}}}{\sqrt{13120+2496\,x-64\,{x}^{2}-24\,{x}^{3}+{x}^{4}}}},$ (52)
$\Psi(x,t=6,c_{13}=3/4)=$
$\sqrt{2}\,{\frac{\sqrt{\sqrt{320+64\,x-8\,{x}^{3}+{x}^{4}}-12+\left(-2+x\right)^{2}}}{\sqrt{320+64\,x-8\,{x}^{3}+{x}^{4}}}},$ (53)
$\Psi(x,t=10,c_{13}=3/4)=$
$\sqrt{2}{\frac{\sqrt{\sqrt{94016+11840\,x-192\,{x}^{2}-40\,{x}^{3}+{x}^{4}}-396+\left(-10+x\right)^{2}}}{\sqrt{94016+11840\,x-192\,{x}^{2}-40\,{x}^{3}+{x}^{4}}}},$ (54)
@@
$\frac{\sqrt{2}}{2}\,{\frac{\sqrt{\sqrt{17-8\,x+18\,{x}^{2}-8\,{x}^{3}+{x}^{4}}-3+\left(-2+x\right)^{2}}}{\sqrt{17-8\,x+18\,{x}^{2}-8\,{x}^{3}+{x}^{4}}}},$ (56)
@@
$\frac{\sqrt{2}}{2}\,{\frac{\sqrt{\sqrt{145-24\,x+146\,{x}^{2}-24\,{x}^{3}+{x}^{4}}-35+\left(-6+x\right)^{2}}}{\sqrt{145-24\,x+146\,{x}^{2}-24\,{x}^{3}+{x}^{4}}}},$ (57)
@@
$\frac{\sqrt{2}}{2}\,{\frac{\sqrt{\sqrt{401-40\,x+402\,{x}^{2}-40\,{x}^{3}+{x}^{4}}-99+\left(-10+x\right)^{2}}}{\sqrt{401-40\,x+402\,{x}^{2}-40\,{x}^{3}+{x}^{4}}}}.$ (58)
$\Psi(x,t=0,c_{13}=-3/4)$ $=$ ${\frac{\sqrt{8+14\,{x}^{2}}\sqrt{2}}{4+7\,{x}^{2}}},$ (59)
$\Psi(x,t=2,c_{13}=-3/4)=$
${\frac{\sqrt{\sqrt{512-896\,x+1008\,{x}^{2}-392\,{x}^{3}+49\,{x}^{4}}-12+7\,\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{512-896\,x+1008\,{x}^{2}-392\,{x}^{3}+49\,{x}^{4}}}},$ (60)
$\Psi(x,t=6,c_{13}=-3/4)=$
${\frac{\sqrt{\sqrt{14848-18816\,x+8624\,{x}^{2}-1176\,{x}^{3}+49\,{x}^{4}}-140+7\,\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{14848-18816\,x+8624\,{x}^{2}-1176\,{x}^{3}+49\,{x}^{4}}}},$
$\Psi(x,t=10,c_{13}=-3/4)=$
${\frac{\sqrt{\sqrt{98816-85120\,x+23856\,{x}^{2}-1960\,{x}^{3}+49\,{x}^{4}}-396+7\,\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{98816-85120\,x+23856\,{x}^{2}-1960\,{x}^{3}+49\,{x}^{4}}}}.$
See these examples in Figures 2 to 7.
@@ 7 Comparison of $\bf\Psi(r,t)$ and $\bf H(r,t)$ in EA and GR
$\Psi(r,t)=\left[{C_{13}}\,{J_{0}\left(wr\right)}+{C_{14}}\,{Y_{0}\left(wr\right)}\right]\cos\left(wt\right).$ (71)
@@
$\Psi(r,t)={J_{0}\left(wr\right)}\cos\left(wt\right).$ (72)
@@
$H(r,t)$ $=$ $-\frac{1}{2}\,wr\left\{2\,\cos\left(wt\right)^{2}J_{0}\left(wr\right)J_{1}\left(wr\right)-\right.$ (73)
$\left.wr\left[J_{0}\left(wr\right)^{2}+J_{1}\left(wr\right)^{2}\right]\right\}.$
See Figure 8 for $\Psi(w)$ and $H(w)$.
@@
Substituting $\cos\left(wt\right)^{2}=[1+\cos\left(2wt\right)]/2$ and $\cos\left(2wt\right)=[\exp\left(2iwt\right)+\exp\left(-2iwt\right)]/2$ into equation (73), we can integrate it and obtain it in terms of the Meijer G function [36], thus
@@
Figure 8: Wave components $\Psi(w)$ and $H(w)$ using $r=1$ and $t=1$ in equations (72) and (73).

### 26 Feb 2024 v2

@@ 1 Introduction
Immediately after completing the theory of General Relativity [1] and proposing the three classical tests [2] to verify the theory experimentally, in 1916, Einstein predicted the existence of gravitational waves as a natural consequence of dynamical spacetime [3]. In this first article [3] on gravitational waves, Einstein concluded that accelerated massive objects generate changes in the spacetime curvature, which propagate like waves with the speed of light. However, he concluded that gravitational waves carry no energy because of a calculational error, which he corrected in 1918 [4]. He also derived a famous formula for the energy loss of a system emitting gravitational radiation. Nevertheless, Einstein tried to disprove the existence of gravitational waves in 1937 [5]. The issue of gravitational waves remained controversial until Bondi [6] settled it for good, which was later explored in a series of papers by Bondi and his collaborators. The first experimental evidence, albeit indirect, came in 1979 with the discovery of the orbital decay of the newly discovered binary pulsar by Hulse and Taylor [7]. Finally, the gravitational waves were directly detected in 2015 by the LIGO and Virgo collaborations [8]. Now, with this direct detection of gravitational waves, we have a new window into the universe and can test our understanding of gravity at higher energies like never before. A fascinating account of the history of gravitational waves can be found in references [9, 10, 11].
@@ 4 Gravitational Wave Equation
Here, we have checked the possible existence of a bifurcation of solutions. We have two possibilities from equation (23): (i) $c_{2}=c_{13}=0$ or (ii) $c_{2}=-c_{13}$. In general we can assume that $H_{0}\neq\Psi_{0}$. In the case (i) we obtain from the equation (16) and from equation (17) the quantity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(-c_{4}+c_{3})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{3}=c_{4}$ into equations (15)-(20) we obtain exactly the GR field equations. In the case (ii) we obtain again from the equation (16) and from equation (17) the quantity $H_{0}$ that must be identical to each other in order to have system consistency, i.e., $\Psi_{0}r(-1+2\Psi_{1}r)(c_{14})(H_{1}-\Psi_{1})=0$. Again, assuming in general that $H_{1}\neq\Psi_{1}$ and substituting the condition $c_{14}=0$ into equations (15)-(20) we obtain exactly the EA field equations. These two cases suggest no GR limit for EA wave equations.
@@ 5 Solution of the Gravitational Wave Equation
${k}={\mathrm{w}}\,\sqrt{1-{c_{13}}},$ (29)
where ${\mathrm{w}}$ is the frequency of the gravitational wave. Notice that in EA, the velocity of the wave can be infinity when $c_{13}\rightarrow 1$, while in GR the velocity of the wave is the speed of the light. We also must have $c_{13}<1$ in order to have a wave.
@@
$\Psi{(r,t)}=f\left(r\right)\cos\left({\mathrm{w}}\,t\right),$ (30)
@@
Let us suppose that $\sqrt{C_{12}}=C_{5}=\omega$, $C_{8}=C_{6}$, $C_{11}=C_{4}$, $C_{10}=C_{3}$, for the sake of simplicity, and using equation (29), we can rewrite the metric functions as
$\Psi{(r,t)}={C_{1}}\,\cos\left({\mathrm{w}}\,t\right)\,J_{0}\left(\sqrt{1-{c_{13}}}\,{\mathrm{w}}\,r\right),$ (42)
$H(r,t)$ $=$ $2{C_{8}}\,J_{0}\left(\omega\,r\right)\left[{C_{3}}\,\cos\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,t\right)+{C_{4}}\,\sin\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,t\right)\right],$
@@
$H(r,t)$ $=$ $\cos\left(\frac{\sqrt{c_{14}}}{\sqrt{{c_{13}}+{c_{2}}}}\,\omega\,\,t\right)J_{0}\left(\omega\,r\right).$ (44)
Comparing equations (42) and (44), we can see that they have the same functional dependence in the coordinates $r$ and $t$, but with different frequencies in both coordinates, ${\mathrm{w}}$ and $\omega$. Section 6 will show that this temporal behavior is different in GR.
@@ 6 Gravitational Wave Pulse with Width
Let us now calculate the superposition of waves of the type of (42) and (44) with amplitude factor as $\exp(-\beta\,{\mathrm{w}})$. Here, we will follow the same procedure used in reference [34]. Thus, we have
@@
assuming that $\beta=1$, $t=0$ or $t=2$ or $t=6$ or $t=10$ and $c_{13}=0$ or $c_{13}=\pm 1/2$, we can write that
$\Psi(x,t=0,c_{13}=1/2)$ $=$ $\,{\frac{\sqrt{2}}{\sqrt{2+{x}^{2}}}},$ (49)
$\Psi(x,t=2,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{100-12\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}-12+2\,\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{100-12\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}}},$ (50)
$\Psi(x,t=6,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{5476-140\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}-140+2\,\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{5476-140\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}}},$ (51)
$\Psi(x,t=10,c_{13}=1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{40804-396\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}-396+2\,\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{40804-396\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}}},$
@@
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{25-6\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}-3+\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{25-6\,\left(-2+x\right)^{2}+\left(-2+x\right)^{4}}}},$ (54)
@@
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{1369-70\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}-35+\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{1369-70\,\left(-6+x\right)^{2}+\left(-6+x\right)^{4}}}},$ (55)
@@
$\frac{1}{2}\,{\frac{\sqrt{\sqrt{10201-198\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}-99+\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{10201-198\,\left(-10+x\right)^{2}+\left(-10+x\right)^{4}}}}.$ (56)
$\Psi(x,t=0,c_{13}=-1/2)$ $=$ ${\frac{\sqrt{2}}{\sqrt{2+3\,{x}^{2}}}},$ (57)
$\Psi(x,t=2,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{100-36\,\left(-2+x\right)^{2}+9\,\left(-2+x\right)^{4}}-12+6\,\left(-2+x\right)^{2}}\sqrt{2}}{\sqrt{100-36\,\left(-2+x\right)^{2}+9\,\left(-2+x\right)^{4}}}},$ (58)
$\Psi(x,t=6,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{5476-420\,\left(-6+x\right)^{2}+9\,\left(-6+x\right)^{4}}-140+6\,\left(-6+x\right)^{2}}\sqrt{2}}{\sqrt{5476-420\,\left(-6+x\right)^{2}+9\,\left(-6+x\right)^{4}}}},$
$\Psi(x,t=10,c_{13}=-1/2)=$
$\frac{1}{2}\,{\frac{\sqrt{2\,\sqrt{40804-1188\,\left(-10+x\right)^{2}+9\,\left(-10+x\right)^{4}}-396+6\,\left(-10+x\right)^{2}}\sqrt{2}}{\sqrt{40804-1188\,\left(-10+x\right)^{2}+9\,\left(-10+x\right)^{4}}}}.$
See these examples in Figures 2 to 7, using the parameters of the Table 1.
Table 1: Summary of the Aether Parameters
Case $c_{13}$ $c_{14}$ $c_{2}$
$A$1/2 2 7/2
$B$ 0 1/2 1/2
$C$ 1/2 3/2 1/4
These parameters must obey the equation (41) and the solar system tests, thus, we must also have $20\leq c_{14}<$ 2, $c_{13}<1$ and $02+c_{13}+2c_{2}>$ 0 [36].
Figure 1: Wave components $\Psi({\mathrm{w}})$ and $H(\omega)$ using $r=1$, $t=1$ and the parameters of Table 1 in equations (42) and (44).
Figure 2: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 3: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and the parameters of Table 1.
Figure 4: Wave component pulse $\Psi$ of width $\beta$, using $C_{1}=1$ in equation (42) and $\alpha=t$, $\beta=1$, $\gamma=\sqrt{1-c_{13}}\,r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 5: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
Figure 6: Wave component pulse $H$ of width $\beta$, using equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$, and the parameters of Table 1.
Figure 7: Wave component pulse $H$ of width $\beta$, using in equation (44) and $\alpha={t\,\sqrt{c_{14}}}/{\sqrt{{c_{13}}+{c_{2}}}}$, $\beta=1$, $\gamma=r$ in equation (46), for four values of $t=0$, $t=2$, $t=6$ and $t=10$ and the parameters of Table 1.
@@ 7 Comparison of $\bf\Psi(r,t)$ and $\bf H(r,t)$ in EA and GR
$\Psi(r,t)=\left[{C_{13}}\,{J_{0}\left({\mathrm{w}}r\right)}+{C_{14}}\,{Y_{0}\left({\mathrm{w}}r\right)}\right]\cos\left({\mathrm{w}}t\right).$ (69)
@@
$\Psi(r,t)={J_{0}\left({\mathrm{w}}r\right)}\cos\left({\mathrm{w}}t\right).$ (70)
@@
$H(r,t)$ $=$ $-\frac{1}{2}\,{\mathrm{w}}r\left\{2\,\cos\left({\mathrm{w}}t\right)^{2}J_{0}\left({\mathrm{w}}r\right)J_{1}\left({\mathrm{w}}r\right)-\right.$ (71)
$\left.{\mathrm{w}}r\left[J_{0}\left({\mathrm{w}}r\right)^{2}+J_{1}\left({\mathrm{w}}r\right)^{2}\right]\right\}.$
See Figure 8 for $\Psi({\mathrm{w}})$ and $H(\omega)$.
@@
Substituting $\cos\left({\mathrm{w}}t\right)^{2}=[1+\cos\left(2{\mathrm{w}}t\right)]/2$ and $\cos\left(2{\mathrm{w}}t\right)=[\exp\left(2i{\mathrm{w}}t\right)+\exp\left(-2i{\mathrm{w}}t\right)]/2$ into equation (71), we can integrate it and obtain it in terms of the Meijer G function [37], thus
@@
Figure 8: Wave components $\Psi({\mathrm{w}})$ and $H(\omega)$ using $r=1$ and $t=1$ in equations (70) and (71).